### v3

# Anamorphic catadioptric telescope and generalized spherical aberration

Abstract
Compensation of the generalized spherical aberration components is discussed for the plane-symmetric and anamorphic optical systems. An example compensation rules are derived for the double-plane symmetric optical system containing reflective and refractive subsystems, which focus light in the orthogonal directions at the image plane. An economical anamorphic extension of the Cassegrain-type cylindrical mirror system has been simulated with a ray-tracing code. Anamorphic telescopes of the introduced type provide independent focal distances in the two orthogonal planes allowing to set up a favorable field of view.
keywords:
anamorphic optical system, spherical aberration, catadioptric system, Cassegrain telescope, cylindrical mirror, toroidal lens
[…]
1 Introduction
Anamorphic optical systems (OS), which possess a double-plane symmetry, have been investigated for many years [1, 2, 3, 4, 5, 6, 7] mainly due to their rich capabilities for image formation. However, they reveal also copious structure of aberrations [1, 8, 9, 10, 4, 11, 12, 13, 14] that remain a subject of research. In particular, S. Yuan and J. Sasian have argued that, in the paraxial region, an anamorphic system can be treated as two associated rotationally symmetric OS, and derived for pure mirror systems the generalized ray-tracing equations that include the aberration terms up to the third order [11, 12]. Recently, set of the third-order aberration coefficients in the paraxial approximation was obtained for the systems of cylindrical mirrors or/and lenses directly from the calculation of the optical path difference without approximating these OS by the pairs of rotationally symmetric OS, and as a particular example a cylindrical analog of the Cassegrain telescope was considered [14]. The present research is devoted to its possible anamorphic extensions that allow for two large magnifications in the orthogonal directions at the image plane and good control of the field of view.
We start from discussing the monochromatic third-order aberrations for either anamorphic or plane-symmetric OS, concentrating on the generalized spherical one, in Sect. 2. Next, we introduce a concept of anamorphic catadioptric telescope with two independent focal distances in Sect. 3. We impose compensation of the spherical aberration in the XZ and YZ planes in its initial design and check the performance using a ray-tracing software. We discuss the results and conclude in 4.
2 Plane symmetries and generalized spherical aberrations
The transverse aberrations for paraxial rays propagating in the $Z$ direction through a generic system of cylindrical optical surfaces, which is symmetric under reflection in the $YZ$ plane and translationally invariant along the $Y$ axis, were obtained in Ref. [14] by extension of the classical methods based on direct calculation of the optical path difference [15]. We consider generalization of this system to a single-plane or double-plane symmetric OS. In the later case the $Z$ axis coincides with the crossing line of the two symmetry planes, and the OS may contain anamorphic surfaces, e.g., toroidal ones. The wave aberration function $W$ being the optical path difference between real and intended ideal wavefronts can be expressed in terms of the four transverse coordinates {$x_{0}$, $y_{0}$, $x$, $y$} of the points $P_{0}(x_{0},y_{0},0)$ and $P(x,y,z)$, in which arbitrary ray crosses the object plane and the exit pupil plane, respectively [16]. The both planes are orthogonal to the $Z$ axis. Symmetry in either $XZ$ or $YZ$ plane as well as double-plane symmetry in both these planes leaves the following six second order invariants for the considered four variables
$x_{0}^{2},\quad y_{0}^{2},\quad x^{2},\quad y^{2},\quad x_{0}x,\quad y_{0}y.$ (1)
The fourth-order wave aberration function is a generic quadratic form in these invariants. The number of its terms can be calculated using the relation between the 2-combinations with repetitions and the binomial coefficient as follows
$\left(\begin{pmatrix}6\\ 2\end{pmatrix}\right)=\begin{pmatrix}6+2-1\\ 2\end{pmatrix}=\frac{7\times 6}{2}=21.$ (2)
This form contains three constant piston terms (involving only $x_{0}$ and $y_{0}$) and two pairs of identical terms with either $x_{0}^{2}x^{2}$ or $y_{0}^{2}y^{2}$. This leaves in the fourth-order wave aberration function
$W^{(4)}\!$ $=a_{1}x^{4}+a_{2}y^{4}+a_{3}x^{2}y^{2}$
$+a_{4}x_{0}x^{3}+a_{5}y_{0}x^{2}y+a_{6}x_{0}xy^{2}+a_{7}y_{0}y^{3}$
$+a_{8}x_{0}^{2}x^{2}+a_{9}y_{0}^{2}y^{2}+a_{10}y_{0}^{2}x^{2}+a_{11}x_{0}^{2}y^{2}+a_{12}x_{0}y_{0}xy$
$+a_{13}x_{0}^{3}x+a_{14}y_{0}^{3}y+a_{15}x_{0}y_{0}^{2}x+a_{16}x_{0}^{2}y_{0}y$ (3)
sixteen different terms that generate third-order aberrations of spherical, comatic, astigmatic (plus field curvature) and distortion types, corresponding to the lines in Eq. (2[17, 18, 19]. For paraxial rays the components of the third-order transverse aberrations can be calculated as
$\delta_{1x}=-\frac{D}{n}\frac{\partial W^{(4)}}{\partial x},\qquad\delta_{1y}=-\frac{D}{n}\frac{\partial W^{(4)}}{\partial y},$ (4)
where $D$ is the distance between the exit pupil plane and the image plane, and $n$ is the refractive index of the material behind the exit pupil. In particular, the generalized spherical (GS) aberrations are given by
$n\delta_{1x}^{GS}$ $=$ $(\tilde{a}_{1}x^{2}+\tilde{a}_{3}y^{2})x,$ (5)
$n\delta_{1y}^{GS}$ $=$ $(\tilde{a}_{3}x^{2}+\tilde{a}_{2}y^{2})y,$ (6)
where $\tilde{a}_{1,2}=4a_{1,2}D$, $\tilde{a}_{3}=2a_{3}D$. Clearly, in the case of a single-plane or double-plane symmetry GS aberrations depend on three parameters, which is similar to the case of a single-plane symmetry combined with translational invariance [14]. The GS aberrations degenerate to the ordinary spherical aberration for $\tilde{a}_{1}=\tilde{a}_{2}=\tilde{a}_{3}$. Hence, for the rays in the $XZ$ or $YZ$ planes the coefficients $\tilde{a}_{1}$ and $\tilde{a}_{2}$, respectively, play the role of a usual spherical aberration coefficient.
We notice that in the case of no symmetry constraints the wave aberration function can contain generically five fourth-order power products of the variables $x$ and $y$: $x^{4}$, $x^{3}y$, $x^{2}y^{2}$, $xy^{3}$ and $y^{4}$. Therefore five parameters can enter GS aberrations.
In the polar coordinates of the exit pupil,
$x=\rho\cos\theta,\qquad y=\rho\sin\theta,$ (7)
for the rays in the $XZ$ and $YZ$ planes the angle $\theta$ is fixed and equal to $k\pi$ and $\pi/2+k\pi$, respectively, where $k$ is an integer. Then conditions for GS aberration compensation in these planes nullify the coefficients $\tilde{a}_{1}$ and $\tilde{a}_{2}$. The remaining coefficient $\tilde{a}_{3}$ enters the quadrifolium (clover) component of GS aberrations:
$n\delta_{1x}^{GS}$ $=a\sin^{2}\theta\cos\theta=\frac{a}{2}\sin 2\theta\sin\theta,$ (8)
$n\delta_{1y}^{GS}$ $=a\cos^{2}\theta\sin\theta=\frac{a}{2}\sin 2\theta\cos\theta,$ (9)
where $a=\tilde{a}_{3}\rho^{3}$, as shown in Fig. 1 that is drawn with Mathematica [20].
Figure 1: Generalized spherical aberrations for $a=1$ and the following chosen values for the parameter set $\tilde{a}_{1}\rho^{3}$, $\tilde{a}_{2}\rho^{3}$} in Eqs. (5)-(7). Left: {1, 1} (solid), {1, 0} (dashed), {0, 1} (dotted) and {0, 0} (dot-dashed). Right: $\tilde{a}_{1}\rho^{3}$, $\tilde{a}_{2}\rho^{3}$} set: {0.5, 1} (solid), {0.5, 0} (dashed), {0, 1} (dotted) and {0, 0} (dot-dashed).
3 GS aberrations for anamorphic catadioptric telescope
In this section we present an anamorphic telescope [21] with a separate focusing along the orthogonal directions $X$ and $Y$ in case of the optical axis coinciding with the $Z$ axis. To be more specific, we discuss catadioptric schemes consisting of the two sets of optical elements: reflective set (I) and refractive set (II). The set (I) contains cylindrical mirror(s) and is responsible for focusing of the collimated light parallel to the Z axis to a focal segment (in case of absence of the second set) at the focal distance $F_{1}$. The set (II) with smaller focal distance $F_{2}$ contains anamorphic lenses and is responsible for the light pencil compressing in the direction of the focal segment and its focusing to a focal point. The set (II), being an anamorphic extension for the set (I), breaks the translational invariance of the OS to the second plane symmetry.
For definiteness we take the translationally invariant set (I) of Cassegrain type [14], which consists of the two cylindrical mirrors: concave primary M1 and convex secondary M2, as shown in Fig. 2. The focal segment $AB$ of the primary M1 is parallel to the focal segment $A_{2}B_{2}$ of the secondary M2. For the set (II) we concentrate on two economical examples, containing either two or three optical surfaces, which are introduced in the following.
Figure 2: The discussed anamorphic OS with a single-lens set (II) in the projections of $XZ$ (left) and $YZ$ (right).
3.1 Single lens set (II)
We add a toroidal lens L drawn in Fig. 3 behind the mirror set (I) in such a way that its symmetry axis coincides with the focal line $A^{\prime}B^{\prime}$ of the set (I). This lens is a body of rotation of its YZ profile with respect to the axis $A^{\prime}B^{\prime}$. We choose L with a biconvex YZ profile. The OS focal point $C$ is in the center of the segment $A^{\prime}B^{\prime}$. The entire anamorphic OS contains only four optical surfaces, indexed by 1 and 2 for M1 and M2, respectively, and 3 and 4 for the surfaces of L. We describe the cylindrical surfaces by the equation
$z=\pm\left[\frac{x^{2}}{2R_{i}}+(1+\kappa_{i})\frac{x^{4}}{8R_{i}^{3}}\right],$ (10)
where overall sign determines the direction of surface convexity, $R_{i}$ is the $i$-th surface’s radius of curvature (always positive) and its only asphericity is in conic constant $\kappa_{i}$.
Figure 3: A schematic isometric view of the optical surfaces 3 (upper) and 4 (lower) of the toroidal lens L.
In the following we discuss aberrations for the considered anamorphic OS, concentrating on GS aberration for brevity. We introduce the conditions for its compensation in the $XZ$ and $YZ$ planes, which correspond to the dotted and dashed lines in Fig. 1, respectively, where the dot-dashed line corresponds to the fulfillment of the both conditions at once. We remark that the introduced method can be applied also to reduction of comatic and other aberration types.
3.1.1 $XZ$ plane
For paraxial rays in the $XZ$ plane and the marginal ray height at the primary the contribution to GS aberration for the system is determined by the function [14]
$A_{3s}^{(1)}=A_{311}+A_{312}\left(\frac{x_{2}}{x_{1}}\right)^{4}\!\!+A_{313}\left(\frac{x_{3}}{x_{1}}\right)^{4}\!\!+A_{314}\left(\frac{x_{4}}{x_{1}}\right)^{4}\!,$ (11)
where $x_{i}$ is the marginal ray height at the $i$-th surface. The aberration coefficients for the mirrors are given by
$A_{311}=-\frac{\kappa_{1}+1}{4R_{1}^{3}},\quad A_{312}=\frac{1}{4R_{2}^{3}}\left[\kappa_{2}+\left(\frac{\mu+1}{\mu-1}\right)^{2}\right],$ (12)
where $\mu>0$ is the transverse magnification of the secondary.
To derive the aberration coefficients $A_{313}$ and $A_{314}$ we replace the refractive surfaces 3 and 4 by the cylindrical ones, which are tangential to them in the $XZ$ plane and are described by Eq. (10) with the overall plus sign. Then we use the following generic expression [14]
$A_{31i}=-\,\frac{1}{8}\!\left[\frac{\kappa}{R^{3}}(n_{1}-n_{0})+\frac{n_{0}}{s_{0}}\left(1+\frac{n_{0}}{n_{1}}\frac{s_{0}}{s_{1}}\right)\!\left(\frac{1}{s_{0}}+\frac{1}{R}\right)^{2}\right]_{i}.$ (13)
The $XZ$ projections for the 3rd and 4th surfaces are the circle arcs with the radii $R_{3}$ and $R_{4}$, respectively, which are centered in the OS focal point $C$, see Fig. 2 (left). We take the refractive index to be $n$ inside the lens and unity outside. For $\kappa_{3x}=\kappa_{4x}=0$, where the index $x$ refers to the section by the $XZ$ plane, we obtain
$A_{313}$ $=$ $-\,\frac{1}{8s_{03}}\left(1+\frac{s_{03}}{ns_{13}}\right)\left(\frac{1}{s_{03}}+\frac{1}{R_{3}}\right)^{2},$ (14)
$A_{314}$ $=$ $-\,\frac{n}{8s_{04}}\left(1+\frac{ns_{04}}{s_{14}}\right)\left(\frac{1}{s_{04}}+\frac{1}{R_{4}}\right)^{2},$ (15)
where $s_{0i}$ and $s_{1i}$ are the lengths of the $XZ$ projections of the chief ray segments connecting the $i$-th surface with the respective object and image point. Using the rule of ‘negative distance’ between a surface and a virtual object (converging incoming beam), we substitute
$s_{0i}=-s_{1i}=-R_{i}$ (16)
with $i=3,4$ in Eqs. (14) and (15), so the coefficients $A_{313}$ and $A_{314}$ vanish. Then condition of compensation of GS aberration in the $XZ$ plane coincides with the one for isolated set (I) [14]:
$F_{XZ}\equiv\kappa_{1}+1-\frac{k_{x}^{4}}{\rho^{3}}\left[\kappa_{2}+\left(\frac{\mu+1}{\mu-1}\right)^{2}\right]=0,$ (17)
where $\rho=R_{2}/R_{1}$ and $k_{x}=x_{2}/x_{1}$. In particular, for a telescope of classical Cassegrain type: $\kappa_{1}=-1$ (paraboloidal M1) and $\kappa_{2}=-(\mu+1)^{2}(\mu-1)^{-2}<-1$ (hyperboloidal M2).
3.1.2 $YZ$ plane
The GS aberration component in the $YZ$ plane for the marginal ray height at the primary can be determined by
$B_{3s}^{(1)}=\tilde{B}_{31}+\tilde{B}_{32}\left(\frac{y_{2}}{y_{1}}\right)^{4}\!\!+B_{33}^{*}\left(\frac{y_{3}}{y_{1}}\right)^{4}\!\!+B_{34}^{*}\left(\frac{y_{4}}{y_{1}}\right)^{4}\!,$ (18)
where $y_{i}$ is the marginal ray distance from the $Z$ axis at the $i$-th surface. The aberration coefficients $\tilde{B}_{3i}=A_{32i}$ vanish. To obtain the coefficients $B_{33}^{*}$ and $B_{34}^{*}$ we use the cylindrical surfaces, which are tangential to the surfaces 3 and 4 in the $YZ$ plane and translationally invariant along the $X$ axis, in contrast to the translation along the $Y$ axis in Ref. [14]. To indicate this difference we replace squiggles with stars. This follows in the interchanged roles of $A\leftrightarrow B$ and $x_{i}\leftrightarrow y_{i}$, which results in $B_{3i}^{*}=A_{31i}$ [14]. By labeling curvature radii in the YZ plane with small letters, according to the proper overall sign in Eq. (10), we obtain
$B_{33}^{*}=-\frac{\kappa_{3}(n-1)}{8r_{3}^{3}}-\frac{1}{8s_{03y}}\left(1+\frac{s_{03y}}{ns_{13y}}\right)\!\left(\frac{1}{s_{03y}}+\frac{1}{r_{3}}\right)^{2}\!,$ (19)
$B_{34}^{*}=-\frac{\kappa_{4}(n-1)}{8r_{4}^{3}}-\frac{n}{8s_{04y}}\left(1+\frac{ns_{04y}}{s_{14y}}\right)\!\left(\frac{1}{s_{04y}}-\frac{1}{r_{4}}\right)^{2}\!,$ (20)
where $r_{i}\equiv r_{iy}$ and $\kappa_{i}\equiv\kappa_{iy}$ refer to the surface that is tangential to the $i$-th surface, and the quantities relevant to the $YZ$ plane are indexed by $y$.
Paraxial equations for refraction on the discussed surfaces tangential to 3 and 4 read
$\frac{n}{s_{13y}}-\frac{1}{s_{03y}}=\frac{n-1}{r_{3}},\qquad\frac{1}{s_{14y}}-\frac{n}{s_{04y}}=\frac{n-1}{r_{4}},$ (21)
where we substitute $s_{03y}=\infty$ and $s_{04y}=s_{13y}-d_{L}$ ($\kappa_{i}=0$ in the paraxial approximation). The thickness $d_{L}$ of of the lens L along the optical axis is given by
$d_{L}=R_{3}-R_{4}.$ (22)
It can not be smaller than the sum of sags $d_{3}$ and $d_{4}$, which can be expressed as
$d_{i}(r_{i},L_{y})=r_{i}\left[1-\sqrt{1-\left(\frac{L_{y}}{2r_{i}}\right)^{2}}\right],\ \ i=3,4,$ (23)
where $L_{y}=AB$ is the lens size along the $Y$ axis (in Fig. 2 $d_{L}=d_{3}+d_{4}$). Using Eq. (21), Eqs. (19) and (20) can be rewritten as
$B_{33}^{*}$ $=-\,\frac{n-1}{8r_{3}^{3}}\left(\kappa_{3}+\frac{1}{n^{2}}\right),$ (24)
$B_{34}^{*}$ $=-\,\frac{n-1}{8r_{4}^{3}}\left[\kappa_{4}+\frac{n}{\nu^{3}}\left(1+n^{2}+n\nu\right)\right]\left(1-n+\nu\right)^{2}$ (25)
with the new dimensionless function
$\nu=n\,\frac{r_{3}}{r_{4}}-(n-1)\,\frac{d_{L}}{r_{4}}.$ (26)
Taking into account that
$\frac{y_{4}}{y_{3}}=1-\frac{n+1}{n}\,\frac{d_{L}}{r_{3}}\approx 1,$ (27)
for $r_{3}=r_{4}\equiv r$ Eq. (18) results in the following condition for compensation of GS aberrations in the $YZ$ plane
$\kappa_{3}+\kappa_{4}=-\frac{1}{n^{2}}-\frac{n}{\nu^{3}}\left(1+n^{2}+n\nu\right)\left(1-n+\nu\right)^{2},$ (28)
where $\nu=n-(n-1)d_{L}r^{-1}$. In Fig. 4 we show this sum of the conic constants as a function of the ratio $d_{L}/r$ for several choices of $n$. Clearly, both $\kappa_{3}$ and $\kappa_{4}$ cannot be zero simultaneously for any reasonable set of parameters, i.e. for not impractically thick lens or high refractive index. Hence, at least one of the lens surfaces must be ‘aspheric’ in its YZ profile to satisfy Eq. (28).
Figure 4: The sum of the conic constants for the surfaces 3 and 4 in the YZ plane versus the ratio of the lens thickness $d_{L}$ to the radius of curvature $r\equiv r_{3}=r_{4}$ for the chosen values of the refractive index $n$: 1.5, 1.7 and 1.9.
3.2 Doublet lens set (II)
Since two refractive surfaces with vanishing conic constants is not enough, we consider the set (II) containing three optical surfaces. Specifically, we choose an anamorphic extension of a refractive doublet objective, contact type, in its standard form with positive front lens and negative rare one, i.e. we take two connected toroidal lenses, positive and negative, with circular YZ profiles. We denote the refractive indices of the positive and negative components by $n$ and $\tilde{n}$, respectively, and chose $n<\tilde{n}$. The respective condition for compensation of GS aberrations reads
$\sum_{i=3}^{5}B^{*}_{3i}\left(\frac{y_{i}}{y_{3}}\right)^{4}=0$ (29)
with
$B_{33}^{*}$ $=-\,\frac{n-1}{8n^{2}r_{3}^{3}},$ (30)
$B_{34}^{*}$ $=-\,\frac{n-1}{8r_{4}^{3}}\,\frac{n}{\nu^{3}}\left[1+\frac{n}{\tilde{n}^{2}}\left(n-\frac{\tilde{n}-n}{n-1}\,\nu\right)\right]\left(1-n+\nu\right)^{2},$ (31)
$B_{35}^{*}$ $=-\,\frac{\tilde{n}(\tilde{n}^{2}+1)}{8(s_{14}-\tilde{d}_{L})^{3}},$ (32)
where we took $r_{5}=\infty$ (plano-concave negative component) for more concise expression of $B_{35}^{*}$, $\nu$ is given by Eq. (26), $d_{L}$ and $\tilde{d}_{L}$ are the thicknesses of positive and negative components of the doublet lens, respectively, and
$s_{14}=\frac{\tilde{n}\nu r_{4}}{n(n-1)-(\tilde{n}-n)\nu}.$ (33)
The distance to the focus behind the doublet lens is given by
$s_{15}=\frac{s_{14}-\tilde{d}_{L}}{\tilde{n}}.$ (34)
The Eq. (29) nearly vanishes for a broad range of values of the parameters.
We remark that the used method can be applied to more complex refractive sets and extended to compensate for other aberrations besides the spherical ones.
3.3 Proof of concept by image simulation
To validate the proposed concept design we consider f/10 anamorphic telescope with two-mirror set (I) and doublet lens set (II), which has two equal focal ratios:
$\frac{F_{1}}{L_{x}}=\frac{F_{2}}{L_{y}}=10,$ (35)
where $L_{x}$ and $L_{y}$ are the sides of its rectangular aperture. To be more specific, we choose the set (I), which is a cylindrical analog of RC-type Cassegrain. For a set (II) we take toroidal analog of the primary spherical aberration and coma free apochromatic doublet [22], since the discussed above economical three-surface contact doublet generates several times larger RMS spot.
We simulated this OS and performed ray-tracing using the optical design software CODE V [23]. The simulation is performed for a on-axis $0^{\circ}$ field point and the wavelength 587.6 nm. The set (I) parameter values we take as follows: $F_{1}=1000$ mm, $R_{1}=450$ mm, $R_{2}=210$ mm, $\kappa_{1}=-1.057$ and $\kappa_{2}=-2.839$. Adopting the apochromatic doublet specification [22] to provide the ratio of foci of $F_{1}/F_{2}\approx 5$, for the set (II) we obtain the following set of curvature radii with the sign factors {$95.75$ mm, $-41.98$ mm, $-41.98$ mm, $-367.74$ mm} and the lens separation of 0.688 mm, where the positive and negative components widths are $d_{L}=3.704$ mm and $\tilde{d}_{L}=2$ mm, and their glass types are FPL52 (OHARA) and ZKN7 (SCHOTT), respectively. The XZ, YZ and prospective views of the anamorphic telescope are shown in Fig. 5.
Figure 5: From top to bottom: XZ view, YZ view and prospective view for the considered realization of an anamorphic catadioptric telescope with the RC-type Cassegrain set (I) and the apochromatic doublet set (II).
Figure 6: Spot diagrams for the discussed anamorphic telescope at the image plane (left) and best focus (right).
The spot diagrams for a on-axis $0^{\circ}$ field point are shown in Fig. 6, where the circle represents a contour of Airy disk, which size is $2.44\times\lambda\times 10$ with 10 standing for the focal ratio. The RMS spot size is 1.5$\pm$0.1 $\mu$m at the image plane and 1.25$\pm$0.1 $\mu$m at the best focus. Clearly, the shown spot diagrams possess double-plane symmetry just like the OS itself.
Next we compare performances of the axially symmetric analogs of isolated sets (I) and (II). The prospective view and spot diagram for such analog of set (I) with the quadratic aperture $L_{x}\times L_{x}$ is shown in Fig. 7. Its RMS spot size is only 0.06$\pm$0.1 $\mu$m. The spot diagrams for the analog of set (II) with the quadratic aperture $L_{y}\times L_{y}$ is shown in Fig. 8. Its RMS spot size is 3.3$\pm$0.1 $\mu$m at the image plane and 1.19$\pm$0.1 $\mu$m at the best focus. For plotting spot diagrams we used the grid ray pattern with 20 rays across the aperture. It is clear from the spot diagrams in Figs. 7 (right) and 8 that aberrations of the considered anamorphic catadioptric telescope are dominated by the aberrations of its refractive set (II).
Figure 7: The prospective view (left) and spot diagram (right) for the axially symmetric analog of isolated set (I) with the aperture 100 mm $\times$ 100 mm.
Figure 8: Spot diagrams at the image plane (left) and best focus (right) for the axially symmetric analog of isolated set (II) with the aperture 20 mm $\times$ 20 mm.
3.4 Field of view of the anamorphic telescope
Let’s compare field of view (FOV) of an anamorphic telescope (AT) possessing the two focal distances $F_{1}$ and $F_{2}$, where $F_{1}>F_{2}$, and FOV of usual rotationally symmetric telescope (RST) with the focal distance $F_{1}$. Suppose that they are equipped with the identical eyepieces, which are specified by the focal distance $f$ and quadratic field stop, i.e. FOV $\varphi\times\varphi$. Then FOV of RST is equal angle:
$\Phi\times\Phi=\frac{\varphi f}{F_{1}}\times\frac{\varphi f}{F_{1}},$ (36)
while FOV of AT is $F_{1}/F_{2}$ times wider in the direction of the focal segment of the set (I):
$\Phi_{1}\times\Phi_{2}=\frac{\varphi f}{F_{1}}\times\frac{\varphi f}{F_{2}}.$ (37)
We conclude that AT possesses $F_{1}/F_{2}$ times wider FOV and uses its CCD camera $F_{1}/F_{2}$ times more effectively with respect to RST of the same rectangular aperture (assuming that CCD camera is a standard quadratic photosensor array).
4 Discussion and conclusion
Following a brief intro on anamorphic OS aberrations and their possible partial compensations, we demonstrated its application to design of an anamorphic telescope, in particular, for GS aberration compensation in the OS symmetry planes. We proposed an anamorphic catadioptric telescope that combines a translationally invariant reflective part and an anamorphic refractive part, which has a rotational symmetry with respect to the axis that is parallel to the direction of translational symmetry of the reflective part and passes through the OS focal point. We outlined derivation of analytical expressions for the GS aberration compensations discussed, considering OS with cylindrical version of the Cassegrain reflector as the reflective part and two examples of the refractive part, which are relatively simple for analytical consideration. Then we simulated the anamorphic telescope combining the cylindrical RC-type Cassegrain with toroidal apochromatic doublet, using the ray-tracing software CODE V [23]. This simulation demonstrated good OS performance, which can be further improved by using more advanced refractive set and introducing corrective lenses.
In general, a wide class of cylindrical versions of known telescopic objectives, either reflective or refractive, can serve as a set (I) and variety of toroidal versions of known axially symmetric refractive telescopic OS can serve as a set (II). (We remark that reflective anamorphic set (II) is also possible [DZ, work in progress].) One of the main limitations comes from avoiding light blocking by the set (II), e.g., the distance to the focus behind the secondary mirror must be sufficiently greater than the focal distance $F_{2}$ of the set (II). The expressions for aberration compensation for the OS symmetry planes XZ and YZ (in the discussed coordinates) are identical to the ones for axially symmetric analogs of the sets (I) and (II), respectively. New analytical expressions can be found for aberration compensation beyond these planes, which is an important direction of future research.
To conclude, we presented a viable concept of an anamorphic telescope of a new type that can possess two large magnifications in the orthogonal directions at the image plane corresponding to the two independent focal distances, which allow to control both dimensions of the filed of view.
Disclosures
The author declares no conflicts of interest.
Code, Data, and Materials Availability
All data in support of the findings of this paper are available within the article.
Acknowledgments
The author thanks Barbara Pintaske and Daniela Ponce for their attention when providing the optical software of the SYNOPSYS.